(2x-1)^2+(x)^2=(2x+1)^2
Answer by tanimachatterjee(60)

(2x-1)^2+(x)^2=(2x+1)^2
or 4x^2 -4x +1 +x^2=4x^2 +4x +1
or x^2 -8x =0
or x(x-8)=0
so either x=0 or x-8=0
x=8